OpenStudy (anonymous):

Choose the point which is a solution to the system of inequalities: y –x + 7 y < 3x + 11

OpenStudy (anonymous):

x=7-y from first equation y<3(7-y)+11 y<-3y+32 4y=32, y=32/4=8 and x can be solved using x=7-y so x=7-8=-1 therefore the solution to this set is at the point (-1,8)

OpenStudy (anonymous):

You didn't post the points. But do it just like in the other problem you posted. Plug the points into the inequalities and see when they're both true.